# ORBIT: An Object Property Reasoning Benchmark for Visual Inference Tasks

### v1

@@ Abstract
While vision-language models (VLMs) have made remarkable progress on many popular visual question answering (VQA) benchmarks, it remains unclear whether they abstract and reason over depicted objects. Inspired by human object categorisation, object property reasoning involves identifying and recognising low-level details and higher-level abstractions. While current VQA benchmarks consider a limited set of object property attributes like size, they typically blend perception and reasoning, and lack representativeness in terms of reasoning and image categories. To this end, we introduce a systematic evaluation framework with images of three representative types, three reasoning levels of increasing complexity, and four object property dimensions driven by prior work on commonsense reasoning. We develop a procedure to instantiate this benchmark into ORBIT, a multi-level reasoning VQA benchmark for object properties comprising 360 images paired with a total of 1,080 count-based questions. Experiments with 12 state-of-the-art VLMs in zero-shot settings reveal significant limitations compared to humans, with the best-performing model only reaching 40% accuracy. VLMs struggle particularly with realistic (photographic) images, counterfactual reasoning about physical and functional properties, and higher counts. ORBIT points to the need to develop methods for scalable benchmarking, generalize annotation guidelines, and explore additional reasoning VLMs. We make the ORBIT benchmark and the experimental code available to support such endeavors.
Codehttps://github.com/AbhishekKolari/ORBIT
Introduction
Object representation and reasoning is one of the core systems of human cognition (Spelke and Kinzler 2007). By abstracting object properties and interactions, humans, including infants with limited real-world knowledge and experience, represent their environment and perform inference (Lipton and Spelke 2004). Object representation and reasoning in humans are driven by spatio-temporal principles of cohesion (objects move as connected and bounded wholes), continuity (objects move on connected, unobstructed paths), and contact (objects do not interact at a distance). Humans perceive object boundaries, represent object shapes that move (partially) out of view, and predict object movement (Aguiar and Baillargeon 1999; Leslie and Keeble 1987). Object perception and inference is not merely descriptive; it involves a multi-stage process from low-level recognition, through categorization, to identification and reasoning (Fields 2016). For instance, to count how many mammals are visible in Figure 1-top, humans recognize the image objects, decide which of them fit the constraint (type) of a mammal, and count the ones that satisfy this criterion.
Figure 1: Visual question answering in ORBIT, with questions about different object properties of varying reasoning complexity on three types of images from different domains.
Category VQAv2 CLEVR TallyQA SCIENCEQA GQA C-VQA CFMM ORBIT
Counting Question
Object Properties Physical
Taxonomic
Functional
@@ Introduction
Image Types Real
@@
Reasoning Complexity Direct Recognition
@@
Table 1: ORBIT versus VQAv2 (Goyal et al. 2017), CLEVR (Johnson et al. 2017), TallyQA (Acharya, Kafle, and Kanan 2018), SCIENCEQA (Lu et al. 2022), GQA (Hudson and Manning 2019), C-VQA (Zhang et al. 2024b), and CFMM (Li et al. 2024).
@@
Recent work on the connection between perception and reasoning has used visual question answering (VQA) in natural language as a diagnostic task for vision-language models (VLMs). Prior work has developed VQA benchmarks for recognition of basic visual elements such as physical attributes or taxonomic category membership (Yuan et al. 2021; Tong et al. 2024), inference of relational properties (Lu et al. 2022; Hudson and Manning 2019), and counterfactual reasoning (Frohberg and Binder 2021; Yu et al. 2023; Li et al. 2024). As objects are central components of images, solving these benchmarks relies on various physical and abstract object properties  (Johnson et al. 2017; Hudson and Manning 2019; Acharya, Kafle, and Kanan 2018).
@@
Current benchmarks exhibit two key limitations. The first limitation is the lack of a delineation between the separate subtasks that comprise object perception and reasoning, namely, object recognition, inference, and counterfactual reasoning. While some works have identified small object sizes and peripheral locations as causes of failure in VLMs (Zhang et al. 2025, 2024a), the lack of separation between these subtasks prevents comprehensive diagnostics of the VLMs’ strengths and weaknesses. Second, current benchmarks are often limited in representativeness of the attributes, as most benchmarks are limited to basic features such as shape and color. While Jiang et al. (2024) recently proposed a systematic benchmark for abstract visual reasoning, it relies on manipulating simple shapes in synthetic environments with minimal noise. Meanwhile, the ability of VLMs to perceive and reason over real-world objects and across image types remains underexplored.
@@
We address this gap in the understanding of the capabilities of current models with a structured and comprehensive evaluation of the ability of VLMs to perceive and reason about a range of property relations of objects using count-based questions in three image types. We make three contributions, which are illustrated in Figure 1: 1. A systematic evaluation frameworkfor object property reasoning, consisting of three image types, three levels of reasoning complexity, and four dimensions of commonsense knowledge. 2. A semi-automaticprocedure for dataset creation and the ORBIT VQA benchmark with 1,080 high-quality counting questions about 360 images. 3. An extensive experimental analysisof a wide range of state-of-the-art VLMs on ORBIT, revealing current models’ limitations compared to humans, especially for photographic images, counterfactual reasoning, functional knowledge, and higher counts.
Related Work
@@
Object Properties in Visual Question Answering. VQA has long served as a diagnostic task for evaluating the visual and reasoning capabilities of AI systems. Early efforts (Antol et al. 2015; Goyal et al. 2017) emphasized general visual recognition of objects, colors, and basic attributes, while subsequent benchmarks (Singh et al. 2019; Mathew, Karatzas, and Jawahar 2021; Lu et al. 2022) targeted domain-specific context. Although many benchmarks across domains (Johnson et al. 2017; Hudson and Manning 2019; Acharya, Kafle, and Kanan 2018) require reasoning over object properties, none of them provide a systematic separation of object property dimensions. Compared to previous VQA benchmarks, ORBIT covers four distinct object dimensions based on commonsense resources (Speer, Chin, and Havasi 2017; Ilievski et al. 2021; Kurtz and Silliman 2021): physical, taxonomic, functional, and relational, and enables fine-grained assessment of object-level reasoning as visualized in Table 1.
@@
Reasoning Complexity in Visual Understanding. Human cognition is inherently compositional, and integrates multiple scene aspects into higher-level reasoning, with complexity progressing from direct recognition to property-level inference, and counterfactual reasoning (Hoffman and Richards 1987; Xu, Huang, and Liu 2021). Initial direct recognition is limited to recognition of basic visual elements such as physical attributes or taxonomic category membership (Yuan et al. 2021; Tong et al. 2024). Inference builds on recognition, targets higher-level functional or relational properties, and requires multi-step abstraction beyond surface-level features (Lu et al. 2022; Hudson and Manning 2019). Counterfactual reasoning involves reasoning about hypothetical changes made in inputs. This type of reasoning is most challenging as it requires understanding and adapting to hypothetical, altered, or out-of-context scenarios (Frohberg and Binder 2021; Yu et al. 2023; Li et al. 2024). Various benchmarks have been proposed to target different types of reasoning (Table 1), yet none of them provide a systematic evaluation across all types simultaneously. In contrast, ORBIT is the first to encompass all three reasoning types and extends evaluation to three image domains, and thereby enables a more comprehensive assessment of models’ reasoning abilities in visual understanding.
A Framework for Object Property Evaluation
@@
We devise a systematic framework for rigorous evaluation of VLMs, integrating object property dimensions (physical, taxonomic, functional), levels of reasoning complexity (recognition, property inference, counterfactual reasoning), and a variety of image types (photographic, animated, AI-generated). We next detail the three framework components.
Object Property Dimensions
@@
Drawing on prior research (Ilievski et al. 2021; Tandon, de Melo, and Weikum 2017) on representing commonsense knowledge about properties and their relations, our framework distinguishes four object property dimensions relevant for count-based VQA. Each dimension captures a distinct facet of human conceptualization and reasoning about objects, their attributes, and relations in visual scenes.
@@
Physical knowledge refers to the dynamics of physical systems based on observable phenomena and fundamental principles (McCloskey, Washburn, and Felch 1983) (e.g., circular objects in Figure 1). Our framework enhances the list of physical properties in prior work, such as shape, color, and size (Table 1), with more complex attributes like the material (wood, metal), state of the object (liquid, solid), and structural characteristics of the object – including part-whole relationships such as has wheels and has legs. Physical properties, such as object qualities, materials, and part-whole relations, are prevalent in commonsense knowledge representation (Tandon, de Melo, and Weikum 2017; Fleming 2017).
@@
Taxonomic knowledge captures an object’s semantic category or class membership, typically expressed as is-a relations (Ilievski et al. 2021). Examples include broad ontological groupings such as biological (e.g., mammals, reptiles), artifact (e.g., furniture, tools), and food (e.g., fruits, vegetables) categories. Unlike physical attributes, taxonomic properties may not be visible from visual input alone. To determine how many mammals are present in an image (Figure 1), for example, visual identification of species needs to be combined with knowledge of their categorization. The taxonomic dimension aligns with taxonomic semantic systems from cognitive science and domain-specific knowledge representation, which group entities based on shared features and categorical similarity (Mirman, Landrigan, and Britt 2017). Taxonomic groupings are also linguistically supported: for instance, Lobben (2012) showed their universality based on similarities in the noun class system of roughly 500 languages in the Bantu language family.
@@
Functional properties express capabilities and general design aspects of an object. This includes attributes such as utilities and capabilities of an object (e.g., means of transportation in Figure 1), affordances as actions an object receives (e.g., breakable, foldable), and operational dependencies for an object to function (e.g., electricity, battery). This dimension primarily aligns with the utility and causal relations identified in commonsense knowledge frameworks (Speer, Chin, and Havasi 2017; Heindorf et al. 2020; Ilievski et al. 2021). Questions about the utility of objects have been included in prior VQA literature (Table 1): for example, the OK-VQA benchmark (Marino et al. 2019) tests for knowledge about implicit functionality based on utility, e.g., recognizing that a broom is for sweeping.
@@
Relational knowledge captures how objects interact, how they can be grouped, and how they are situated relative to one another within a visual context. Attributes include spatial relationships (e.g., on top of, hanging from) and grouping relations (e.g., couple, flock). Questions like ”How many couples are visible?” or ”How many objects can be seen hanging from the wall?” evaluate a model’s ability to reason about such properties. The relational dimension is motivated by the spatial and relational-other relations in commonsense reasoning (Ilievski et al. 2021). It is limited to spatial and contextual groupings of instances, making it complementary to the internal relations of objects in part-whole relationships in the physical dimension and taxonomic groupings in the taxonomic dimension.
Reasoning Complexity
@@
The four object property dimensions can support questions with various levels of complexity. We introduce three levels of reasoning complexity: direct recognition, property inference, and counterfactual reasoning, exemplified in Figure 1. We expect that the progression from direct recognition to counterfactual reasoning requires an increasingly high level of abstraction to answer a question, thus increasing in difficulty (Xu, Huang, and Liu 2021).
@@
Direct Recognition questions are features that can be detected by observation and general categorical knowledge. Questions at this level can be recognized through an atomic step of perception and reside at the physical and taxonomic property dimensions. Since answers can be derived from specific visual regions or patterns, recognition questions primarily assess object recognition abilities, e.g., the top question in Figure 1 only requires identifying land animals and categorizing them as mammals.
@@
Property Inference questions are targeted at the functional and relational property dimensions, which require deeper abstraction, introducing multi-step reasoning. Due to the properties having codependencies among objects and underlying relations across different image regions, this would push models to generalize beyond surface-level features. For example, in Figure 1, the middle question requires identifying a bicycle, a car, and a plane, categorizing them as vehicles, and then reasoning about their utility in transportation.
@@
Counterfactual Reasoning questions probe the VLM’s ability to reason about hypothetical, altered, or out-of-context scenarios. Counterfactual questions span any of the four property dimensions. In addition to visual recognition and inference, counterfactual questions also require commonsense knowledge and flexible abstraction (e.g., creating a contextualized procedure that identifies remaining circular objects besides half of the clocks in the bottom question of Figure 1), which makes them the most complex of the three reasoning categories in our framework.
Image Types
@@
To evaluate VLM generalization across visual domains, we incorporate a diverse set of image types, recognizing that model performance may vary (and typically worsen) when moving from synthetic to real-world images (Johnson et al. 2017; Hudson and Manning 2019).
@@
Photographic images represent naturalistic settings, featuring complex textures, lighting variations, and cluttered scenes. They are realistic, i.e., all depicted objects are encountered in the real world. Photographs present challenges like occlusions, viewpoint variation, and natural ambiguity, demanding a high level of visual grounding from models.
@@
Animated images are simplified or stylized and contain minimal visual noise, as well as limited levels of detail when compared to photographic images. Humans typically create animated images in a digital format, and their objects may or may not be realistic. Object boundaries in animated images may be easier to detect, but exaggerated representations may pose a novel challenge.
@@
AI-generated images include both photorealistic and stylized images, created by prompting generative models. These images test a model’s robustness to domain shifts, as they may deviate from typical real-world physics, object proportions, and other visual conventions. AI-generated images typically feature simpler compositions and include objects that are implausible in practice, such as half-animals or glasses floating in the air.
@@
Figure 2: ORBIT construction pipeline: image collection, candidate QA formulation, and quality assurance. n and m stand for number of images and questions, respectively.
ORBIT Benchmark Construction

### v2

# A Study of Commonsense Reasoning over Visual Object Properties

@@ Abstract
Inspired by human categorization, object property reasoning involves identifying and recognizing low-level details and higher-level abstractions. While current visual question answering (VQA) studies consider multiple object properties, such as size, they typically blend perception and reasoning and lack representativeness in terms of reasoning and image categories, making it unclear whether and how vision-language models (VLMs) abstract and reason over depicted objects. To this end, we introduce a systematic evaluation framework comprising images of three representative types, three reasoning levels of increasing complexity, and four object property dimensions, informed by prior work on common sense. We develop a procedure to instantiate this framework in two VQA object reasoning benchmarks: OPTICS-CNT, comprising 360 images paired with 1,080 multi-level, count-based questions, and OPTICS-CMP, with 2.1k comparison questions. Experiments with 12 state-of-the-art VLMs in zero-shot settings reveal significant limitations relative to humans, with the best-performing model achieving below 40% counting and 70% comparison accuracy. VLMs struggle particularly with photographic images, counterfactual reasoning, physical and functional properties, and higher counts. We make the OPTICS benchmark data and code available to support future work on scalable benchmarking methods, generalized annotation guidelines, and advanced reasoning VLMs.
1 Introduction
Object representation and reasoning is one of the core systems of human cognition (Spelke and Kinzler, 2007). By abstracting object properties and interactions, humans, including infants with limited real-world knowledge and experience, represent their environment and perform inference (Lipton and Spelke, 2004). Object representation and reasoning in humans are driven by spatio-temporal principles of cohesion (objects move as connected, bounded wholes), continuity (objects move along connected, unobstructed paths), and contact (objects do not interact at a distance). Humans perceive object boundaries, represent object shapes that move (partially) out of view, and predict object movement (Aguiar and Baillargeon, 1999; Leslie and Keeble, 1987). Importantly, object perception and inference is not merely descriptive; it involves a multi-stage process from low-level recognition, through categorization, to identification and reasoning Fields (2016). For instance, to count how many mammals are visible in Figure 1-top, humans recognize the image objects, decide which of them fit the constraint (type) of a mammal, and count those.
Figure 1: VQA in OPTICS-CNT, with questions about different object properties of varying reasoning complexity on three types of images from different domains.
Category VQAv2 CLEVR TallyQA SCIENCEQA GQA C-VQA CFMM OPTICS
Question Counting
Comparison
Physical
Object Taxonomic
Property Functional
@@ 1 Introduction
Image Type Real
@@
Reasoning Direct Recognition
@@
Table 1: Comparison of OPTICS to closely related existing benchmarks: VQAv2 Goyal et al. (2017), CLEVR Johnson et al. (2017), TallyQA Acharya et al. (2018), SCIENCEQA Lu et al. (2022), GQA Hudson and Manning (2019), C-VQA Zhang et al. (2024b), and CFMM Li et al. (2024).
@@
Recent work on the connection between perception and reasoning has used visual question answering (VQA) in natural language as a diagnostic task for vision-language models (VLMs). Existing VQA benchmarks evaluate recognition of basic visual elements such as physical attributes or taxonomic category membership Yuan et al. (2021); Tong et al. (2024), inference of relational properties Lu et al. (2022); Hudson and Manning (2019), and counterfactual reasoning Frohberg and Binder (2021); Yu et al. (2023); Li et al. (2024). As objects are central components of images, solving these benchmarks relies on extracting and reasoning over physical and abstract object properties  Johnson et al. (2017); Hudson and Manning (2019); Acharya et al. (2018).
@@
Current VQA evaluation exhibits two key limitations. The first is the lack of a delineation between subtasks for object perception and reasoning, i.e., object detection, inference, and counterfactual reasoning. While some studies have identified small object sizes and peripheral locations as causes of failure in VLMs (Zhang et al., 2025, 2024a), the lack of subtask separation hinders comprehensive diagnosis of their strengths and weaknesses. Second, we note a limited representativeness of the covered evaluation dimensions, usually focusing on basic features such as shape and color (Goyal et al., 2017; Antol et al., 2015).  Jiang et al. (2024)’s recent systematic evaluation of abstract visual reasoning covers different types, configurations, and shapes. Given its focus on manipulating simple shapes in synthetic environments with minimal noise, the ability of VLMs to perceive and reason over real-world objects across image types remains underexplored.
@@
We address this gap in the understanding of the capabilities of current models with a comprehensive, fine-grained evaluation of the ability of VLMs to perceive and reason given a representative range of object properties, reasoning complexities, and image types. We make three contributions (Figure 1): 1. A systematic evaluation frameworkfor object property reasoning, consisting of three image types, three levels of reasoning complexity, and four dimensions of commonsense knowledge. 2. A semi-automatic procedurefor dataset creation, resulting in two VQA benchmarks based on our framework: OPTICS-CNT with 1.1k high-quality counting questions about 360 images, and OPTICS-CMP with 2.1k comparison questions about image pairs. [^1] OPTICS stands for Object Property Reasoning Tasks for Evaluating Image-based Common Sense. 3. An extensive experimental analysisof state-of-the-art (SotA) VLMs on OPTICS, revealing fine-grained limitations of current models for photographic images, counterfactual reasoning, functional knowledge, and higher counts.
2 Related Work
@@
Object Properties in VQA. VQA has long served as a diagnostic task for evaluating the perceptual and reasoning capabilities of AI systems. Early efforts Antol et al. (2015); Goyal et al. (2017) emphasized general visual recognition of objects, colors, and basic attributes, while subsequent benchmarks Singh et al. (2019); Mathew et al. (2021); Lu et al. (2022) targeted domain-specific context. Although many benchmarks across domains Johnson et al. (2017); Hudson and Manning (2019); Acharya et al. (2018) require reasoning over object properties, none of them provides a systematic separation of object property dimensions. Compared to previous VQA benchmarks, our framework covers four distinct object dimensions based on commonsense resources Speer et al. (2017); Ilievski et al. (2021); Kurtz and Silliman (2021): physical, taxonomic, functional, and relational, and enables fine-grained assessment of object-level reasoning as visualized in Table 1. Moreover, while both counting and comparison questions have been relatively well-covered in prior work, our evaluation enables compositional evaluation by comparing counts across images rather than object properties (e.g., which object is larger Li et al. (2024)).
@@
Visual Reasoning Complexity. Human cognition is inherently compositional and integrates multiple scene aspects into higher-level reasoning, with complexity progressing from direct recognition to property-level inference, and counterfactual reasoning Hoffman and Richards (1987); Xu et al. (2021). Initial direct recognition is limited to recognition of basic visual elements such as physical attributes or taxonomic category membership Yuan et al. (2021); Tong et al. (2024). Inference builds on recognition, targets higher-level functional or relational properties, and requires multi-step abstraction beyond surface-level features Lu et al. (2022); Hudson and Manning (2019). Counterfactual reasoning involves reasoning about hypothetical changes made in inputs. This type of reasoning is most challenging, as it requires understanding and adapting to hypothetical, altered, or out-of-context scenarios Frohberg and Binder (2021); Yu et al. (2023); Li et al. (2024). Benchmarks that target different types of reasoning exist (Table 1), yet none systematically test reasoning levels simultaneously. Inspired by compositional evaluation frameworks in abstract visual reasoning Jiang et al. (2024), OPTICS is the first to encompass all three reasoning types and go beyond previous work to study VLM performance across three image types, thereby enabling a more comprehensive assessment of models’ perception and reasoning abilities. To support fair evaluation, we focus on counting and comparison questions in line with prior benchmarks Acharya et al. (2018); Johnson et al. (2017); Li et al. (2024).
3 A Commonsense Framework for Evaluating Object Property Reasoning
@@
We devise a systematic framework for rigorous evaluation of VLMs, integrating object property dimensions (physical, taxonomic, functional, relational), levels of reasoning complexity (recognition, property inference, counterfactual), and diverse image types (photographic, animated, AI-generated). We next detail its three components.
3.1 Object Property Dimensions
@@
Drawing on prior research Ilievski et al. (2021); Tandon et al. (2017) on representing commonsense knowledge about properties and their relations, our framework distinguishes four object property dimensions relevant for counting and comparison VQA. Each dimension captures a distinct facet of human conceptualization and reasoning about object attributes and relations in visual scenes.
@@
Physical knowledge refers to the dynamics of physical systems based on observable phenomena and fundamental principles (McCloskey et al., 1983) (e.g., circular objects in Figure 1). Physical properties, such as object qualities, materials, and part-whole relations, are prevalent in commonsense knowledge representation (Tandon et al., 2017; Fleming, 2017). Our framework enhances the list of physical properties in prior VQA studies, such as shape, color, and size (Table 1), with more complex attributes like the material (wood, metal), state of the object (liquid, solid), and structural characteristics of the object – including part-whole relationships such as has wheels.
@@
Taxonomic knowledge captures an object’s semantic category or class membership, typically expressed as is-a relations (Ilievski et al., 2021). Examples in our framework include broad ontological groupings, e.g., biological (e.g., mammals, reptiles), artifact (e.g., furniture, tools), and food (e.g., fruits, vegetables) categories. Unlike physical attributes, taxonomic properties may not be visible from visual input alone: to determine how many mammals are present in an image (Figure 1), visual identification of species must be combined with knowledge of their taxonomic classification. The taxonomic dimension aligns with taxonomic semantic systems from cognitive science and domain-specific knowledge representation, which group entities based on shared features and categorical similarity (Mirman et al., 2017).
@@
Functional properties express attributes such as utilities and capabilities of an object (e.g., means of transportation in Figure 1), affordances as actions an object receives (e.g., breakable, foldable), and operational dependencies for an object to function (e.g., electricity, battery). This dimension primarily aligns with the utility and causal relations identified in commonsense knowledge frameworks (Speer et al., 2017; Heindorf et al., 2020; Ilievski et al., 2021). Questions about the utility of objects have been included in VQA studies (Table 1): for example, the OK-VQA benchmark (Marino et al., 2019) tests for knowledge about object utility, e.g., a broom is for sweeping.
@@
Relational knowledge captures how objects interact, how they can be contextually grouped, and how they are situated relative to each other within a visual context. Attributes include spatial relationships (as in "How many objects can be seen hanging from the wall?") and contextual grouping relations (as in "How many couples are visible?"), inspired by commonsense knowledge organization (Ilievski et al., 2021). Its emphasis on relations between multiple object instances complements the focus on properties of single objects in taxonomic and part-whole physical properties.
3.2 Reasoning Complexity
@@
The four object property dimensions can support questions with varying reasoning complexity. Our three levels of reasoning complexity: direct recognition, property inference, and counterfactual reasoning are exemplified in Figure 1. We expect that the progression from direct recognition to counterfactual reasoning requires an increasingly high level of abstraction to answer a question, and a shifting emphasis from perception to reasoning, thus increasing in difficulty (Xu et al., 2021).
@@
Direct Recognition questions ask about features that can be detected by observation and general categorical knowledge. Questions at this level can be answered through an atomic step of perception and reside at the physical and taxonomic property dimensions. For example, the top question in Figure 1 only requires identifying land animals and categorizing them as mammals.
@@
Property Inference questions target the functional and relational property dimensions, which require deeper abstraction, introducing multi-step reasoning. It studies the ability of models to generalize beyond surface-level features by focusing on codependencies among objects and underlying relations across different image regions. For example, in Figure 1, the middle question requires identifying a bicycle, a car, and a plane, categorizing them as vehicles, and then reasoning about their utility in transportation.
@@
Counterfactual Reasoning questions probe the ability to reason about hypothetical, altered, or out-of-context scenarios. Counterfactual questions span any of the four property dimensions. In addition to visual recognition and inference, counterfactual questions also require commonsense knowledge and flexible abstraction (e.g., creating a contextualized procedure that identifies remaining circular objects besides half of the clocks in the bottom question of Figure 1), thus forming the most complex reasoning level in our framework.
3.3 Image Types
@@
To study VLM generalization across visual domains, we cover various image types, expecting that accuracy varies (often worsens) when moving from synthetic to real-world images (Johnson et al., 2017; Hudson and Manning, 2019).
@@
Photographic images represent naturalistic settings, featuring complex textures, lighting variations, and cluttered scenes. They are realistic, i.e., all depicted objects are found in the real world. Photographs present challenges such as occlusions, viewpoint variation, and natural ambiguity, requiring an effective visual grounding.
@@
Animated images are simplified or stylized and contain minimal visual noise, as well as limited levels of detail when compared to photographic images. Humans typically create animated images in digital form, and their objects are not necessarily realistic. Object boundaries in animated images may be easier to detect, but exaggerated representations may pose a novel challenge.
@@
AI-generated images include both photorealistic and stylized images, created by prompting generative models. They test robustness to domain shifts by deviating from visual conventions such as real-world physics and object proportions. AI-generated images typically feature simpler compositions and include implausible objects, such as half-animals or floating glasses.
@@
Figure 2: OPTICS-CNT and OPTICS-CMP construction pipeline: image collection, candidate QA formulation, quality assurance, and pairwise combination. n and m are number of images and QAs, respectively.
4 Benchmark Construction